# Observable CP-violation in charmed baryons decays with SU(3) symmetry analysis

### v3

@@ Abstract
Figure 1: Topology diagrams for the charmed baryon decays into an octet baryon and a light meson. Here $T_{1-7}$ show the Cabibbo-favored channels $c\to u\bar{d}s$. $T_{8,9}$ describe the new effects in the Cabibbo-suppressed $c\to u\bar{d}d(\bar{s}s)$ induced by $\textbf{f}_{b}$ and $\textbf{g}_{b}$ respectively.
Note that the diagrams $T_{1}-T_{2}$ and $T_{4}-T_{5}$ have the same topological structures, where one selects $\mathcal{M}_{T_{1}}=\mathcal{M}_{T_{2}}$ and $\mathcal{M}_{T_{4}}=\mathcal{M}_{T_{5}}$. The number of independent topological diagrams is reduced from seven to five, consistent with the latest TDA analysis in Ref. Zhong:2024qqs . As demonstrated in Ref. He:2018joe , each topological diagram in Fig. 1 corresponds to more than one TDA amplitude. Therefore, it is assumed that the absolute value of the TDA amplitude with the same topological diagram should be equal. By defining the form factor $\mathcal{M}_{T_{i}}=G_{F}\bar{u}(f_{T_{i}}-g_{T_{i}}\gamma_{5})u,\;i=1\sim 7$, one obtains the relations as
$|A_{1,15}|=A_{T_{1}}=|A_{2,16}|=A_{T_{2}},\quad|A_{3,5,9}|=A_{T_{3}},$
$|A_{4,6,10}|=A_{T_{4}}=|A_{7,8,11}|=A_{T_{5}},$
@@
The assumptions are approximately supported by the numerical results within $1\sigma$ in Table. 1 and it shows the diagram $T_{1,2}$ actually give the main contribution. However, some form factors contradict the assumption, such as $|f_{1,15}|-|f_{2,16}|\sim 0.06$, $|f_{T_{17,18}}|-|f_{19}|\sim 0.02$ and $|g_{17,19}|-|g_{18}|\sim 0.05$. To focus on the largest effects, we ignore the discrepancy involving $f_{19}$ and primarily consider $|f_{1,15}|>|f_{2,16}|$ and $|g_{17,19}|>|g_{18}|$ in the subsequent analysis. This implies that there are other effects in the topology diagrams. For exploring the possible new effects, one defines the form factors $\textbf{f}_{b}$ and $\textbf{g}_{b}$ to absorb these effect as $\textbf{f}_{b}=f_{T_{1}}-f_{T_{2}}$ and $\textbf{g}_{b}=-g_{18}-g_{T_{7}}$. The corresponding topology diagrams $T_{8,9}$ are given in Fig. 1. Then the IRA form factors are expressed by TDA form factors and $\textbf{f}_{b},\textbf{g}_{b}$ as
@@
$g^{b}_{6}=-3g_{T_{1}}-2g_{T_{4}},\quad g^{b}_{15}=-g^{e}_{15}=\frac{\textbf{g}_{b}}{2}+g_{T_{6}}+g_{T_{7}},$
$g^{c}_{6}=g_{T_{6}},\quad g^{d}_{6}=3g_{T_{1}}-g_{T_{7}}+\frac{\textbf{g}_{b}}{2},\quad g^{d}_{15}=g_{T_{6}}.$ (5)
@@
The above analysis demonstrates that the fits using the IRA method are largely consistent with the TDA assumption. Subsequently, we determine the strong phase from the perspective of topological diagrams. Since the strong phase arises from potential contributions of intermediate on-shell states in the decay process PDG , the imaginary part of the form factor in Eq. 4 corresponding to each topological diagram is incorporated into our global analysis. Thus, the imaginary part of the IRA form factor in our fitting is derived using Eq. Observable CP-violation in charmed baryons decays with SU(3) symmetry analysis under the assumption: ${\mathcal{I}}m(\textbf{g}_{b})=0$. Additionally, we assume $\mathcal{I}m(f_{T_{3}})=0$ due to the global phase.
According to Eq.Observable CP-violation in charmed baryons decays with SU(3) symmetry analysis, the IRA form factor $f^{b}_{6}$ depends solely on the new effect, potentially leads to significant CP violation. To study the CPV, the possible sources of the new form factor $\textbf{f}_{b}$ are discussed. Since the form factor arises from the difference between topological diagrams $T_{1}$ and $T_{2}$ in Fig.1, the most likely source is the SU(3) symmetry breaking effect. Especially when long-distance interactions enter the interaction vertex in Fig.1, the flavor symmetry breaking effect must be involved. Additionally, the penguin operator in Eq. 1 can also contribute to the new form factor $\textbf{f}_{b}$. When only flavor symmetry is considered, the difference in the Lorentz structure at the Hamiltonian vertex can be neglected. Consequently, the penguin operator has the same Hamiltonian matrix structure with the $O_{1,2}$ in TDA. However, in IRA, the penguin operator only contributes to $H_{3}$, which is neglected in our work. Therefore, the new form factor $\textbf{f}_{b}$ introduced by the conflicting parts of IRA and TDA must be considered.
According to the above analysis, the form factor can be divided into two parts $\textbf{f}_{b}=\textbf{f}^{S}_{b}+\textbf{f}^{P}_{b}e^{i\phi^{P}}$, where $\textbf{f}^{S}_{b}$ comes from the SU(3) symmetry breaking and $\textbf{f}^{P}_{b}e^{i\phi^{P}}$ represents the penguin operator contribution with its weak phase $\phi^{P}=-1.147\pm 0.026$. Including the two part of $\textbf{f}_{b}$, the form factors in our global analysis are defined as
@@
By leveraging the equivalence between IRA and TDA methods, our global analysis extends to complex form factors. By considering the SU(3) symmetry breaking and penguin operator as the source of form factor $\textbf{f}_{b}$, the corresponding contributions of SU(3) symmetry breaking and penguin operator are determined through global analysis respectively. Then we further explore CP violation in anti-triplet charmed baryon two-body decay processes. Our analysis yields non-zero $A_{CP}$ in Eq. 7. Considering its error, the observable CP violation can be observed at the order of $O(10^{-2})$. We strongly advocate for measuring these processes.

### v4

@@ Abstract
Figure 1: Topology diagrams for the charmed baryon two body decays. Here $T_{1-7}$ corresponding to the TDA amplitude in Eq.Observable CP-violation in charmed baryons decays with SU(3) symmetry analysis. $T_{8,9}$ describe the new effects induced by $\textbf{f}_{b}$, $\textbf{g}_{b}$ and $\textbf{g}_{c}$ respectively.
Note that diagrams $T_{1}-T_{2}$ and $T_{4}-T_{5}$ share the similar topological structures, and their differences arise from operators $O_{1}$ and $O_{2}$. As demonstrated in Ref. He:2018joe , each topological diagram in Fig. 1 corresponds to more than one TDA amplitude. Therefore, it is assumed that the absolute value of the TDA amplitude with the same topological diagram should be equal. By defining the form factor $\mathcal{M}_{T_{i}}=G_{F}\bar{u}(f_{T_{i}}-g_{T_{i}}\gamma_{5})u,\;i=1\sim 7$, one obtains the relations as
$|A_{1,15}|=A_{T_{1}},\quad|A_{2,16}|=A_{T_{2}},\quad|A_{3,5,9}|=A_{T_{3}},$
$|A_{4,6,10}|=A_{T_{4}},\quad|A_{7,8,11}|=A_{T_{5}},$
@@
[…]
For exploring the possible new effects, one defines the form factors $\textbf{f}_{b}$ and $\textbf{g}_{b,c}$ to absorb these effect as as $\textbf{f}_{b}=f_{T_{1}}-f_{T_{2}}$, $\textbf{g}_{b}=g_{T_{1}}-g_{T_{2}}$ and $\textbf{g}_{c}=-g_{18}-g_{T_{7}}$. The corresponding topology diagrams $T_{8,9}$ are given in Fig. 1. Then the IRA form factors are expressed by TDA form factors and $\textbf{f}_{b},\textbf{g}_{b,c}$ as
@@
$g^{b}_{6}=-3g_{T_{1}}+\frac{3}{2}\textbf{g}_{b}-2g_{T_{4}},\quad g^{b}_{15}=\frac{\textbf{g}_{c}}{2}+g_{T_{6}}+g_{T_{7}},$
$g^{c}_{6}=g_{T_{6}},\quad g^{d}_{6}=3g_{T_{1}}-\frac{3}{2}\textbf{g}_{b}-g_{T_{7}}+\frac{\textbf{g}_{c}}{2},\quad g^{d}_{15}=g_{T_{6}},$
$g^{e}_{15}=-g_{T_{6}}-g_{T_{7}}-\frac{\textbf{g}_{c}}{2}-\frac{3}{2}\textbf{g}_{b}.$ (5)
@@
The above analysis demonstrates that the fits using the IRA method are largely consistent with the TDA assumption. Subsequently, we determine the strong phase from the perspective of topological diagrams. Since the strong phase arises from potential contributions of intermediate on-shell states in the decay process PDG , the imaginary part of the form factor in Eq. 4 corresponding to each topological diagram is incorporated into our global analysis. We assume $\mathcal{I}m(f_{T_{3}})=0$ due to the global phase.
[…]
Based on the potential origins of these new form factors, their contributions can be estimated by referring to previous studies. The contribution of the penguin operator is estimated to be less than 10% Wang:2024ztg ; Xing:2024nvg . Although there are many indications suggesting SU(3) symmetry-breaking effects in charmed baryon decays, the previous global fit still indicates that the symmetry-breaking effect can be considered negligible within its error He:2021qnc . Regarding the FSR, recent studies indicate that the strong phase in the parity-conserving form factor corresponding to FSR is nearly zero He:2024pxh . Therefore, in our work, we can neglect the imaginary component of the parity-conserving new form factor as ${\mathcal{I}}m(\textbf{g}_{b/c})=0$, since $\textbf{g}_{b/c}$ appears in conjunction with other TDA form factors. Although the contribution of these new form factors may be small, we find that the new form factor $\textbf{f}_{b}$ contributes exclusively to the IRA form factor $f_{6}^{b}$, suggesting that the contribution of $\textbf{f}_{b}$ cannot be neglected.
According to the above analysis, the new form factor can be expressed as $\textbf{f}_{b}=\textbf{f}^{S}_{b}+\textbf{f}^{P}_{b}e^{i\phi^{P}}$, where $\textbf{f}^{S}_{b}$ arises from SU(3) symmetry breaking and non-factorizable QCD contributions, which do not generate a weak phase. The term $\textbf{f}^{P}_{b}e^{i\phi^{P}}$ represents the contribution from the penguin operator and final state radiation (FSR) effects, with the weak phase $\phi^{P}=-1.147\pm 0.026$. Including the two part of $\textbf{f}_{b}$, the form factors in our global analysis are defined as
@@
By leveraging the equivalence between IRA and TDA methods, our global analysis extends to complex form factors. By considering the SU(3) symmetry breaking and penguin operator as the source of form factor $\textbf{f}_{b}$, the corresponding contributions of SU(3) symmetry breaking and penguin operator are determined through global analysis respectively. Then we further explore CP violation in anti-triplet charmed baryon two-body decay processes. Our analysis yields non-zero $A_{CP}$ in Eq. 7. Considering its error, the observable CP violation can be observed up to the order of $O(10^{-2})$. We strongly advocate for measuring these processes.